# NeRF-To-Real Tester: Neural Radiance Fields as Test Image Generators for Vision of Autonomous Systems

### v1

@@ V Experimental Evaluation
  • RQ2: How effective is N2R-Tester at generating diverse images of a scene or object that uncover inconsistent behaviours in an image processing SUT?
  • RQ3: How efficient is N2R-Tester at detecting inconsistencies in the behaviour of an image processing SUT?
Research questions RQ1 and RQ2 address the two core requirements of Sect. III. The spatiotemporal consistency is satisfied by-design, and it is evaluated in NeRF papers.
@@ V-A Training Data
We record five videos of underwater scenes using a manually operated underwater drone equipped with a single front-facing camera. We mimic a trajectory of an AUV during inspections of vertical planar walls. Each video lasts approximately one minute. Images are extracted at 25 FPS and the resolution of $1920\times 1080$. The scenes contain walls covered by various types of marine growth. We estimate the camera poses with colmap [17] (see Tbl. I).
@@
We use two image data sets available in nerfstudio [18] (plane, dozer), that are captured with a handheld camera device. The nerfstudio data sets contain images and estimated camera poses for each image. We record three additional scenes with a handheld camera, containing a car, a truck and a bike and extract images at 25 FPS with image resolution of $1920\times 1080$ pixels. We again use colmap for camera pose estimation. The images of all data sets have the respective vehicle roughly at the image centre and the camera trajectory mimics a UAV circling around the vehicle at an approximately fixed level above ground. See also Tbl. I.
@@
TABLE I: Training data sets. Each of the #total images has an associated camera pose estimation. 90% are used for NeRF training. 10% are used for NeRF in-training evaluation.
@@ V-B Systems Under Test
We evaluate N2R-Tester on four AUV components as SUTs: two DNN-based IPDs as well as two popular state-of-the-art non-DNN IPDs.
  • SuperPoint: SuperPoint, a convolutional DNN, predicts interest points by calculating heatmaps of likely interest point locations. It is trained on a large dataset with partially manual and automatically labelled ground truth data [19].
  • UnSuperPoint: UnSuperPoint is a DNN-based IPD, that incorporates point regression and a mechanism similar to Non-maximum suppression directly in its architecture which helps finding evenly distributed interest points across an image. It is trained in a self-supervised fashion [20].
  • Scale-Invariant Feature Transform (SIFT): SIFT finds interest points in an image by convolving Gaussian kernels at different scales across the image and subsequently selecting local extrema in difference-of-Gaussian pyramids [21].
  • Oriented FAST and Rotated BRIEF (ORB): ORB is based on the FAST point detector, which measures local gradients in pixel intensities in circular patches around each pixel [22].
We use the OpenCV implementations of SIFT and ORB [23]. SuperPoint with pre-trained weights is available on github [24]. UnSuperPoint is proprietary software and currently not publicly available with pre-trained weights. We obtained it from the authors.
@@
We evaluate N2R-Tester on four UAV components as SUTs; two small DNNs with $\sim$4-5M parameters, selected for their capacity to be deployed on modern quadrocopters, as well as two large DNNs. We choose MobileNet [25], EfficientNet [26], XCeption [27] and VGG16 [28], all as implemented in Keras [29]. All models have a convolutional architecture and are trained on ImageNet-1k [30]. Each image is downsampled to a model-specific resolution. We use a recommended downsampling method of OpenCV, namely “inter-area” downsampling. See Tbl. II.
DNN Model Size Input Res. Top-1 Acc. #Params
MobileNet 16 224$\times$224 70.4% 4.3M
EfficientNetB0 29 224$\times$224 77.1% 5.3M
XCeption 88 299$\times$299 79.0% 22.9M
VGG16 528 224$\times$224 71.3% 138.4M
TABLE II: Subject classifier SUTs, data after Keras  [31]. Size is for a file of pre-trained weights in MB. Each DNN uses a specific input layer resolution. The top-1 accuracy refers to the accuracy on the ImageNet-1k validation set.
V-C Performance Metrics for Systems Under Test
@@
We consider a total of four SUT metrics, three for the UAV mission and two for the AUV mission, and inject them into our MT framework:
@@
  • Image Classification: Cosine Similarity $\uparrow$: The final layer of a DNN produces predictions for each class in the data set. As our UAV SUTs are trained on a subset of ImageNet with 1k classes, the output of the DNN is a vector of 1000 scores. The cosine similarity is a measure of similarity between two DNN outputs. It is the cosine angle between the two vectors; two parallel vectors have a cosine similarity of 1, two orthogonal vectors have a cosine similarity of -1.
  • Image Classification: L2 Norm $\downarrow$: The final layer of the DNN produces predictions for each class in the data set. As our UAV SUTs are trained on a popular subset of ImageNet with 1k classes, the output of each DNN is a vector of 1000 scores. The L2 norm between two DNN outputs is the Euclidean squared distance between the two output vectors.
  • Image Classification: Class Invariance $\uparrow$: When comparing two different images and their respective DNN output vectors, we consider whether the highest probability is given to the same class. We define class invariance as the percentage of images in a set where two prediction vectors have the highest probability for the same class.
  • IPD: Repeatability Score $\uparrow$: To report the performance IPDs we use the repeatability score between two sets of interest points. By calculating a homography between two images of the same scene, we can project the first set of interest points onto a second image and find the percentage of matched points [32]. We allow a maximum distance of 2 pixels to consider two points a match.
  • IPD: Interest Point Spread $\uparrow$: As an intuitive measure of the quality of interest points, we calculate their Jaccard Index, i.e. the ratio between the intersection and union of the areas that the interest points cover in an image. A distribution across an image is a desired property of interest points to improve stability [33].
V-D NeRF Models
@@
The structure-from-motion tool colmap selects a minimum 300 suitable images of a captured scene by balancing sufficient visual information and processing speed. Colmap subsequently estimates the camera poses for each image. For our NeRF models we choose a network architecture called Nerfacto, available in nerfstudio [18] because of its balance between training/rendering speed and visual quality. The NeRF training process takes 90% of image/camera-pose combinations to learn each scene by updating its network weights. The remaining 10% are left out for NeRF in-training-evaluation on image fidelity metrics PSNR, SSIM, LPIPS continuously during the training process. We train a total of ten NeRF models on underwater and aerial scenes listed in Tbl. I. We predominantly used the default parameters in nerfstudio to train our models with a fixed number of 30k training steps where each training step comprises 4096 ray samples.
V-E Evaluation for N2R-Tester
@@
In order to demonstrate the efficacy of N2R-Tester in producing realistic and diverse test images, and thus answering RQ1, RQ2, and RQ3, we inject our SUTs, NeRF models and MRs into our framework. Additionally, we employ artificial image mutations suggested by authors of DeepXplore [6] to enable a baseline comparison. Note that in their work, the image mutations are used to generate test images to find DNN test inputs that trigger differential SUT behaviour and cause high neuron coverage. We are not interested the neuron coverage metric since it has shown inconclusive efficacy in other works [34]. We still choose the DeepXplore mutations because they are engineered artificially, thus we are curious to their level of realism and comparative output in SUT behaviour.
(a) Real
(b) NeRF
(c) Real
(d) NeRF
Figure 6: Visual quality of NeRF-generated images: (a) Real images of three underwater scenes of an AUV wall inspection mission. (b) NeRF-generated pose-equivalent images as (a) in the respective NeRF models dory1, dory2 and dory3. (c) Real images of three scenes of an UAV vehicle inspection mission. (d) NeRF-generated pose-equivalent images as (c) in the respective models bike, car and dozer.
VI Results
In this section, we discuss the experiment results of N2R-Tester, and its ability to test our SUTs, i.e. interest point detectors and image classifiers. All test images are rendered at a resolution of (960, 540, 3) pixels.
VI-A RQ1: Realness
Visual quality of NeRF-generated images
Figure 6 shows a range of real and NeRF-generated images of underwater and vehicle scenes to convince the reader of the visual quality of images produced by rendering an image with a NeRF. The images generated by the underwater models dory1, dory2 and dory3 demonstrate an almost indistinguishable resemblance with the real world scene, with a relatively small loss in visual definition of specific features. The images produced by bicycle, car and dozer show a similar visual quality and convince by rendering subtle details with consistency, for example in the bicycle spokes in bicycle. Some short-comings can be seen in the foreground of dozer in the slightly under-defined representation of the street tarmac.
The NeRF reconstruction process is initialised with random pixels, which gradually converge to a geometrically consistent scene during training. During inference this presents an advantage; NeRFs fail gracefully when faced with ambiguity in the reconstruction, i.e. when rendering an image at a view with insufficient radiance samples. This can however result in “artifacts” or “ghost walls”, i.e. image patches appearing as random noise or blurry (Fig. 12). We find that an effective test for detecting ghost walls in a NeRF image is if an IPD returns fewer than the preset number of interest points, in our case 100. Researchers in the NeRF community are in a race trying to outperform each other in mitigating these pixel effects [35]. Nerfacto models assume the real-world scenes to be static, but other types of NeRF architectures aim at modelling dynamic scenes [36, 37]. Additionally, NeRFs assume that the real-world scene contains one main light source, and they tend to struggle in reconstructing surfaces and objects that reflect light such as mirrors and glass surfaces. Figure 12 exemplifies a few of these phenomena. We are curious as to how exactly these phenomena potentially affect the performance of the SUTs, however this exploration marks a whole line of research in its own right. We found the near and far bound parameters $t_{n}$, $t_{f}$ in Eq. 1 supported convergence during the reconstruction process for dory1 and dory3 based on evaluation loss. We report the image metrics PSNR, SSIM and LPIPS for each scene in Fig. 8 in blue. Notice that worse scores on the image metrics for dozer and plane correlate with worse scores on the SUT metrics in Fig. 7. Transformations $\tau$ often present on average equal or higher values for PSNR and LPIPS than DeepXplore mutations, and hence suggest a higher level of realness.
(a) Avg. Cos. Similarity $\uparrow$
(b) Avg. L2 Norm $\downarrow$
(c) Class Invariance $\uparrow$
(d) Avg. Repeatability $\uparrow$
(e) Avg. IP Spread $\uparrow$
SUT Performance under reality-to-NeRF domain shift
We investigate the sensitivity of each SUT to the domain shift from real to NeRF-generated images according to the selected SUT metrics. We consider the set of real images $S_{\textrm{real}}$ and their pose-equivalent counterparts in $S_{\textrm{nerf}}$ (e.g. Fig. 6 first column compared to second column). The significance of the SUT metric values under the domain shift is twofold; it can be seen as 1) a measure of quality in the NeRF reconstruction and realness of NeRF-generated images, and 2) an indication of relative performance of each SUT when presented with visually similar content but small pixel perturbations. See Fig. 7 for average values of each SUT metric in their respective UAV/AUV application. If we passed the exact same image to any SUT, we would expect a value of 1 for cosine similarity, class invariance, repeatability and IP spread, as well as a value of 0 for the L2 Norm. Any deviation of these values can be seen as an indication of the two properties described above. For each UAV scene, MobileNet shows the lowest scores throughout 13/15 settings compared to other SUTs. ORB is the IPD with the most consistent interest points under the domain shift in all AUV scenes according to the repeatability metric. The interest points of SuperPoint and SIFT appear to find interest points in the most consistent areas compared to the other SUTs. The models dozer and plane consistently show an overall lower levels of realness in the reconstruction, considering all SUTs show lower scores for these models in all settings.
We are curious whether the image quality metrics PSNR, SSIM, LPIPS have correlation with the UAV/AUV-associated SUT metrics under the reality-to-NeRF domain shift. We consider the real images $S_{\textrm{real}}$ and their pose-equivalent images in $S_{\textrm{nerf}}$. Intuitively, the better an image $i_{\textrm{nerf}}$ is reconstructed, i.e. has better values for PSNR, SSIM and LPIPS, then the features between pairs of $i_{\textrm{real}}$ and $i_{\textrm{nerf}}$ should also have higher agreement, and therefore higher SUT metrics. Thus the images show a higher level of realness. We find there is significant dependence for 38 out of 48 SUT metric / image quality metric combinations as seen in Tbl. III, based on Spearman rank correlations up until a p-value of 0.05.
Repeatability IP Spread
PSNR SSIM LPIPS PSNR SSIM LPIPS
SIFT 0.45 (1.71e-83) 0.57 (1.18e-144) -0.52 (1.32e-114) 0.15 (4.21e-18) 0.22 (5.48e-37) -0.18 (1.21e-25)
ORB 0.39 (1.37e-61) 0.56 (1.13e-135) -0.47 (1.76e-89) 0.31 (1.01e-73) 0.30 (8.76e-70) -0.38 (1.89e-112)
SuperPoint 0.38 (2.99e-59) 0.74 (2.66e-288) -0.28 (6.20e-31) 0.12 (6.26e-13) 0.22 (2.94e-36) -0.11 (1.11e-10)
UnSuperPoint 0.47 (8.34e-93) 0.70 (4.57e-244) -0.44 (1.05e-77) 0.27 (2.50e-54) 0.20 (9.23e-32) -0.30 (2.80e-71)
Cosine Similarity L2 Norm
MobileNet 0.47 (4.14e-105) 0.46 (5.59e-104) -0.39 (3.19e-69) 0.01 (4.49e-01) 0.02 (3.36e-01) -0.01 (5.76e-01)
EfficientnetB0 0.27 (2.75e-34) 0.41 (3.96e-77) -0.50 (5.99e-123) 0.01 (7.33e-01) 0.00 (9.24e-01) -0.02 (3.06e-01)
XCeption 0.31 (1.86e-45) 0.34 (5.50e-53) -0.37 (4.79e-64) 0.00 (8.29e-01) -0.01 (6.11e-01) -0.01 (7.11e-01)
VGG16 0.52 (1.90e-136) 0.45 (8.61e-97) -0.56 (1.10e-156) -0.00 (8.12e-01) -0.01 (6.50e-01) -0.00 (8.63e-01)
TABLE III: Spearman Rank correlation coefficients and associated p-values (in brackets) between image quality metrics PSNR, SSIM, LPIPS and SUT metrics under the reality-to-NeRF domain shift in pose-equivalent images in $S_{\textrm{real}}$ and $S_{\textrm{nerf}}$.
(a) Avg. PSNR $\uparrow$
(b) Avg. SSIM $\uparrow$
(c) Avg. LPIPS $\downarrow$
Figure 8: Comparison of image metrics per scene under different pose transformations and mutations.
VI-B RQ2: Diversity
Since each NeRF model has its own scale, we design targeted pose transformations for each NeRF model individually. We consider the reality-to-NeRF domain shift a transformation in itself and call it $\tau_{0}$. We additionally perform six types of pose transformations in the NeRF model:
@@ VI-B RQ2: Diversity
We manually design the transformations, such that the main object or viewpoint of the scene is in full view at all times. To achieve this, for example when translating the camera upward, we apply a small rotation in pitch downward. This viewpoint requirement is assessed by manual inspection and takes a matter of seconds. By applying these transformations, we multiply the amount of test images in each scene sixfold as reported in Tbl. IV. Note that the transformations (1, 2), (3, 4) and (5, 6) increasingly deviate from the original camera path and should therefore produce more diverse image samples. We choose the following three image mutations by DeepXplore as our baseline, because from a practical point of view, they come closest to real-life failure modes in cameras:
@@
Number of Inconsistencies
We report the number of inconsistencies per SUT in Fig. 10 and per scene in Fig. 11 according to our MT framework with three different threshold values for $\epsilon=[0.1,0.2,0.5]$. This means that we consider a deviation of a SUT metric by more than 10%, 20%, 50% an inconsistency in the behaviour of a SUT. The strictest test is $\epsilon=0.1$ and should show the highest number of inconsistencies. Considering only NeRF-generated it can be said that transformations further away from the original path such as $\tau_{4},\tau_{5},\tau_{6}$ show a higher number of inconsistencies per SUT for cosine similarity and IP spread. The biggest challenge for SUTs according to L2 norm seems to be the reality-to-NeRF shift in general. There is a difference in number of inconsistencies in DeepXplore mutations $m_{2},m_{3}$ compared to $m_{1}$. This is to be expected, since $m_{1}$ is a simple change in image brightness and should not challenge the SUT behaviour to a great extend. For L2 Norm, DeepXplore mutations $m_{2},m_{3}$ quickly loose inconsistencies as the threshold $\epsilon$ increases. Generally, NeRF-images seem to target different aspects of SUT behaviours, that uncover a different range of inconsistencies than DeepXplore mutations. According to the IP spread metric, on average, IPDs seem more challenged by viewpoint changes of the NeRF-generated images than by mutated images.
Examples of Inconsistent Behaviour
See Fig. 9 for two examples of inconsistent SUT behaviour.
Figure 9: Examples of inconsistent SUT behaviours: Image (a) transformed by $\tau_{4}$ is classified as ImageNet1k class label “trash can” by MobileNet with 0.5285 confidence while having average image metric values (PSNR 15.77, SSIM 0.45, LPIPS 0.72). Image (b) transformed by $\tau_{4}$ has a low repeatability score of 0.37 on SuperPoint. The low score is caused by a low agreement of points from the associated real image (green) compared to points found in this image (red). Image (b) has good metric values PSNR 35.12, SSIM 0.70, LPIPS 0.60.

### v2

@@ V Experimental Evaluation
  • RQ2: How effective is N2R-Tester at generating diverse images of a scene or object that uncover inconsistent behaviors in an image processing SUT?
  • RQ3: How efficient is N2R-Tester at detecting inconsistencies in the behavior of an image processing SUT?
Research questions RQ1 and RQ2 address the two core requirements of Sect. III. The spatiotemporal consistency is satisfied by design, and it is evaluated in NeRF papers.
@@ V-A Training Data
We record five videos of underwater scenes using a manually operated underwater drone equipped with a single front-facing camera. We mimic a trajectory of an AUV during inspections of vertical planar walls. Each video lasts approximately one minute. Images are extracted at 25 FPS and the resolution of $1920\times 1080$. The scenes contain walls covered by various types of marine growth. We estimate camera poses with colmap [11] (Tbl. I).
@@
We use two image data sets available in nerfstudio [12] (plane, dozer), captured with a handheld camera device. We record three additional scenes with a handheld camera, showing a car, a truck and a bike and extract images at 25 FPS with image resolution of $1920\times 1080$ pixels. We use colmap for camera pose estimation. The images of all data sets have the respective vehicle roughly at the image center and the camera trajectory mimics a UAV circling the vehicle at an approximately fixed level above ground. See also Tbl. I for details on the trained NeRF models.
@@
TABLE I: Details of NeRF models: All models are our own except plane, dozer, available in nerfstudio.
@@ V-B Systems Under Test
We evaluate N2R-Tester on four AUV components as SUTs: two DNN-based IPDs as well as two popular state-of-the-art non-DNN IPDs. We use the OpenCV implementations of SIFT and ORB [13]. SuperPoint with pre-trained weights is available on github [14]. UnSuperPoint is proprietary software and currently not publicly available with pre-trained weights.
@@
We evaluate N2R-Tester on four UAV components as SUTs; two small DNNs with $\sim$4-5M parameters, selected for their capacity to be deployed on modern quadrocopters, as well as two large DNNs. We choose MobileNet , EfficientNet, XCeption, and VGG16, all as implemented in Keras [15]. All models have a convolutional architecture and are trained on ImageNet-1k [16].
V-C Image Quality Metrics
We consider the following image quality metrics, when comparing an image to a reference image:
  • Peak-Signal-To-Noise Ratio (PSNR): based on mean squared error, scaled to pixel ranges of the underlying distribution.
  • Structural Similarity Index (SSIM): estimates similarity according to luminance and contrast [17].
  • Learned Perceptual Image Patch Similarity (LPIPS): measures similarity in a DNN feature representation [18].
V-D Performance Metrics for Systems Under Test
@@
We consider five SUT metrics, three for the UAV mission and two for the AUV mission, and inject them into N2R.
@@
  • Image Classification: Cosine Similarity $\uparrow$: cosine angle between two DNN outputs; two parallel vectors have a cosine similarity of $1$, two orthogonal vectors of $-1$. DNN outputs are the probability distributions over 1000 classes.
  • Image Classification: L2 Norm $\downarrow$: Euclidean squared distance between the two DNN output vectors with minimum value of 0 and the maximum value of $\sqrt{2}$.
  • Image Classification: Class Invariance $\uparrow$: equivalence of the class with the highest probability of two DNN outputs.
  • IPD: Repeatability Score $\uparrow$: ratio of matched points by projecting one set of points into a second image with a homography [19], with maximum discrepancy of 2 pixels.
  • IPD: Interest Point Spread $\uparrow$: ratio between the intersection and union of the coverage areas of interest points, demonstrating point stability [20].
V-E NeRF Models
@@
The structure-from-motion tool colmap selects a minimum 300 suitable images of a scene by balancing sufficient visual information and processing speed. Colmap subsequently estimates the camera poses for each image. For our NeRF models we choose a network architecture called Nerfacto, available in nerfstudio [12] because of its balance between training/rendering speed and visual quality. We train a total of ten NeRF models on underwater and aerial scenes listed in Tbl. I. We mainly use the default parameters in nerfstudio with 30k training steps and 4096 ray samples per step.
V-F Evaluation for N2R-Tester
@@
To demonstrate the efficacy of N2R-Tester in producing realistic and diverse test images, and thus answering RQ1, RQ2, and RQ3, we inject our SUTs, NeRF models and MRs into our framework. The MRs between test inputs are the chosen spatial transformations $\tau$. The MRs between test outputs are the changes on SUT performance on metric $q$, that is $q(f(i_{\textrm{nerf}}),f(i_{\tau}))>\epsilon$, with three threshold values for $\epsilon$. Additionally, we employ artificial image mutations suggested by authors of DeepXplore [5] to enable a baseline comparison. Note that in their work, the image mutations are used to generate test images to find DNN test inputs that trigger differential SUT behavior and cause high neuron coverage. We are not interested the neuron coverage metric since it has shown inconclusive efficacy in other works [21]. We still choose the DeepXplore mutations because they are engineered artificially, thus we are curious to their level of realism and comparative output in SUT behavior. Since each NeRF model has its own scale, we design targeted pose transformations for each NeRF model individually. We consider the reality-to-NeRF domain shift a transformation in itself and call it $\tau_{0}$. We perform six types of pose transformations in the NeRF model:
@@
We manually design the transformations, such that the main object or viewpoint of the scene is in full view at all times. Therefore the image classifier should return the same object class at all camera poses. Similarly the IPD should give consistent interest points under our transformations. To achieve this, for example when translating the camera upward, we apply a small rotation in pitch downward. By applying these transformations, we multiply the amount of test images in each scene sixfold. Note that the pairs of transformations $\tau_{1}$/$\tau_{2}$, $\tau_{3}$/$\tau_{4}$) and $\tau_{5}$/$\tau_{6}$ increasingly deviate from the original camera path and should therefore produce more diverse images. We choose three image mutations by DeepXplore as our baseline, because from a practical point of view, they come closest to real-life failure modes in cameras (examples in Fig. 5):
@@
(a) $m_{1}$
(b) $m_{2}$
(c) $m_{3}$
Figure 5: DeepXplore mutations of real images of a dozer, with (a) changes in brightness, (b) a random pixel patch and (c) multiple black patches, imitating camera failure modes.
VI Results
(a) Real image
(b) N2R $\tau_{0}$
(c) Real image
(d) N2R $\tau_{0}$
Figure 6: (a) A real image of an underwater scene of an AUV mission and (b) its pose-equivalent image in NeRF model dory1. (c) A real image of an UAV mission, and (d) its pose-equivalent image in NeRF model bike.
(a) Avg. PSNR $\uparrow$
(b) Avg. SSIM $\uparrow$
(c) Avg. LPIPS $\downarrow$
Figure 7: Image metric values for PSNR, SSIM, and LPIPS of images generated by N2R (red) vs DeepXplore (blue).
In this section, we discuss the experiment results of N2R-Tester, and its ability to test our SUTs, i.e. IPDs and image classifiers. All test images are rendered at a resolution of 960x540 pixels.
VI-A RQ1: Realness
Visual quality of NeRF-generated images
Figure 6 shows real and NeRF-generated images of underwater and vehicle scenes to convince the reader of the visual quality of NeRF-generated images. The image generated by the underwater model dory1 demonstrate an almost indistinguishable resemblance with the real world scene. The image produced by bicycle shows a similar visual quality and convince by subtle details, e.g. the bicycle spokes. We found the near and far bound parameters $t_{n}$, $t_{f}$ in Eq. 1 supported convergence during the reconstruction process for dory1 and dory3 based on evaluation loss. We report the image quality metrics PSNR, SSIM and LPIPS for each scene in Fig. 7. N2R images of the underwater models show better values for PSNR and LPIPS than DeepXplore mutations, suggesting a higher level of realness. Under closer investigation, $\tau_{0}$ and $m_{3}$ show the least amount of visual change, and also show the best image quality metrics, as expected.
SUT Performance under reality-to-NeRF domain shift
We investigate the sensitivity of each SUT to the domain shift from real to NeRF-generated images according to the selected SUT metrics. We consider the set of real images $S_{\textrm{real}}$ and their pose-equivalent counterparts in $S_{\textrm{nerf}}$ (e.g. Fig. 6 first column compared to second column). The significance of the SUT metric values under the domain shift is twofold; it can be seen as 1) a measure of quality in the NeRF reconstruction and realness of NeRF-generated images, and 2) an indication of relative performance of each SUT under the reality-to-NeRF domain shift. See Fig. 8 for average values of each SUT metric in their respective UAV/AUV application. If we passed the exact same image to any SUT, we would expect a value of 1 for cosine similarity, class invariance, repeatability and IP spread, and 0 for the L2 Norm. Any deviation of these values can be seen as an indication of the two properties described above. For each UAV scene, MobileNet shows the lowest scores throughout 13/15 settings compared to other SUTs. ORB is the IPD with the most consistent interest points under the domain shift in all AUV scenes according to the repeatability metric. The interest points of SuperPoint and SIFT appear to find interest points in the most consistent areas compared to the other SUTs. Thus, depending on the vSLAM implementation of choice, we can now make more informed decisions on which IPD or classifier to employ to improve performance downstream tasks with more robustness.
(a) Avg. Cos. Similarity $\uparrow$
(b) Avg. L2 Norm $\downarrow$
(c) Class Invariance $\uparrow$
(d) Avg. Repeatability $\uparrow$
(e) Avg. IP Spread $\uparrow$
We are curious whether the image quality metrics PSNR, SSIM, LPIPS have correlation with the UAV/AUV-associated SUT metrics under the reality-to-NeRF domain shift. We consider the real images $S_{\textrm{real}}$ and their pose-equivalent images in $S_{\textrm{nerf}}$. Intuitively, the better an image $i_{\textrm{nerf}}$ is reconstructed, i.e. has better values for PSNR, SSIM and LPIPS, then the features between pairs of $i_{\textrm{real}}$ and $i_{\textrm{nerf}}$ should also have higher agreement, and therefore higher SUT metrics. We find there is significant dependence for 38 out of 48 SUT metric / image quality metric combinations, based on Spearman rank correlations up until a p-value of 0.05.
VI-B RQ2: Diversity